OpenStudy (anonymous):

evaluate: limit as x approaches negative infinity 3/(e^x-5) my professor said the answer should be negative infinity but it's not

OpenStudy (aum):

Is it \(\large \lim_{x \rightarrow -\infty}\frac{3}{e^x - 5}\) or \(\large \lim_{x \rightarrow -\infty}\frac{3}{e^{x - 5}}\) ?

OpenStudy (anonymous):

the first one

OpenStudy (aum):

\[\large \lim_{x \rightarrow -\infty}\frac{3}{e^x - 5} = -\frac 35\]

OpenStudy (anonymous):

can you briefly explain how you got that answer, please?

OpenStudy (aum):

Because \[\large \lim_{x \rightarrow -\infty}e^x = 0\]

OpenStudy (aum):

But \[\large \lim_{x \rightarrow -\infty}\frac{3}{e^{x - 5}} = \infty \] That is, if (x-5) is an exponent, then the limit is +infinity.

OpenStudy (aum):



OpenStudy (anonymous):

oh okay, thank youu!